OpenStudy (anonymous):

the three angles of a triangle have equal measures. the sum of their measuers is 180n degrees. what is the measure of each angle?

OpenStudy (anonymous):

60 dgrees

OpenStudy (saifoo.khan):

60 obv.

OpenStudy (anonymous):

3x=180 3x/3=180/3 x=60 degrees

OpenStudy (anonymous):

oohhh i didnt understand the wording. thanks!

OpenStudy (anonymous):

you don't have to be rude seafood

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

inside joke

OpenStudy (saifoo.khan):

Nope.

OpenStudy (saifoo.khan):

haha, i wasn't rood.

OpenStudy (anonymous):

i guess not really cuz everyone knows ur a jerk

OpenStudy (anonymous):

@saifoo

OpenStudy (anonymous):

gtg cya :)

OpenStudy (anonymous):

it's ok it is just for now but sooner or later you will gonna understand a word problem much easier just answer and solve though

OpenStudy (saifoo.khan):

Haha, now see who's being rude? :P

OpenStudy (anonymous):

ya im really sik and kind of out of it. thanks soo much!

OpenStudy (saifoo.khan):

lol, whtever. but the answer is 60.

OpenStudy (anonymous):

keep calm balliegirl24

OpenStudy (anonymous):

i dont care. sorry. gtg cya. gosh @saifoo

OpenStudy (anonymous):

Bye! :)